# Quantitative polynomial cohomology and applications to $\textrm{L}^{p}$-measure equivalence

### v1

@@ 4 Applications for virtually nilpotent groups
(Gromov [gromov-polynomial-growth]) A finitely generated group $(\Gamma,S)$ has polynomial growth if and only if $\Gamma$ is virtually nilpotent.
@@
(Mal’cev, Gromov) Let $\Gamma$ be a finitely generated nilpotent group. Then for every $d\in\mathbb{N}_{\geq 2}$, $\Gamma$ is of type $F_{d}$ and the combinatorial filling function $\mathrm{cFV}^{d}_{\Gamma}$ is polynomially bounded.
Finiteness properties of finitely generated nilpotent groups come from the fact that they are cocompact lattices in their Mal’cev completions [malcev]. It is hard to single out one reference for the polynomiality of filling functions of nilpotent groups. The book [varopoulos-potential] allows to see it in different ways. For instance, in the book’s terminology, nilpotent Lie groups are NC [varopoulos-potential, Section 2.2.1], and if a group is NC then it is polynomially retractable [varopoulos-potential, Theorem 7.10] and being polynomially retractable implies having the polynomial filling property [varopoulos-potential, Definition 7.12], where the latter means that the more classical Lipschitz versions of the filling functions $\mathrm{FV}_{\Gamma}^{k}$ are polynomially bounded for every $k\in\mathbb{N}_{\geq 2}$. The Federer-Fleming deformation technique shows that $\mathrm{cFV}_{\Gamma}^{k}\sim\mathrm{FV}_{\Gamma}^{k}$ (see for instance [bader-sauer, Theorem 6.6]).
The third property that we need is the higher analogue of Shalom’s property $H_{T}$ [shalom2004harmonic, p. 125]. To our knowledge, this property was first shown for nilpotent groups by Blanc [blanc].
@@
(Blanc [blanc, 10.5], [guichardet-livre, p. 243]) Let $G$ be a connected and simply connected nilpotent Lie group. Then for every $k\geq 1$ and for every continuous unitary representation $\pi$ without $G$-invariant vectors we have:
@@
By Shapiro’s lemma for cocompact lattices [shalom2004harmonic, 3.2.2], this property holds for torsion-free finitely generated nilpotent groups. By Shapiro’s lemma for finite index subgroups, it also holds for general finitely generated virtually nilpotent groups.
@@ 4.2 Betti numbers of nilpotent groups and $\textrm{L}^{p}$-ME
The next Corollary is obtained by specializing the previous Theorem to the case of the trivial representation. Its proof is essentially the same as [shalom2004harmonic, 4.1.1].
@@
Let $\Lambda$ and $\Gamma$ be finitely generated virtually nilpotent groups. Let $d_{\Gamma}=\inf\{d\in\mathbb{R}_{>0},\#S_{\Gamma}(n)\lesssim n^{d}\},k\in\mathbb{N}_{\geq 2}$. Let $N_{\Lambda,k}=\prod_{i=2}^{k}\deg\mathrm{cFV_{\Lambda}^{i}},N_{\Gamma,k}=\prod_{i=2}^{k}\deg\mathrm{cFV_{\Gamma}^{i}}$ and $N=\max\{N_{\Lambda,k},N_{\Gamma,k}\}$ and let $M>\alpha_{k-1}(N)$ where the functions $\alpha_{i}$ are defined in Theorem C. Suppose that $\Lambda$ and $\Gamma$ are mutually coboundedly $\textrm{L}^{p}$-ME for some $p>d_{\Gamma}k+2M+1$ and let $(\Omega,X_{\Lambda},X_{\Gamma},m)$ be an $\textrm{L}^{p}$-integrable mutually cobounded ME-coupling.
@@ 4.3 Nilpotent Lie algebras and gradings
Every finitely generated virtually nilpotent group contains a finite index subgroup that is torsion-free and nilpotent [raghunathan, Lemma 4.6]. Moreover, every finitely generated, torsion-free, nilpotent group $\Gamma$ arises as a cocompact lattice in a simply connected nilpotent Lie group $G$, called the Mal’cev completion of $\Gamma$, whose corresponding Lie algebra $\mathfrak{g}$ has a basis with rational structural constants [malcev] [raghunathan, 2.18]. One can compute Betti numbers of $\Gamma$ using Betti numbers of the Lie algebra $\mathfrak{g}$ (which are defined using Lie algebra cohomology). Indeed, since $\Gamma$ is cocompact in $G$, we have that
@@
for every $n$ (this is can be seen using [shalom2004harmonic, 3.2.2] and [guichardet-livre, p. 243]). Moreover, we also have $b_{n}(G)=b_{n}(\mathfrak{g})$ [guichardet-livre, Chapitre II], where the latter means Betti numbers in the sense of Lie algebra cohomology.
@@
We recall some results on $\textrm{L}^{p}$-measure equivalence of nilpotent groups. On one hand, Austin showed that $\mathrm{L}^{1}$-ME finitely generated groups have bilipschitz asymptotic cones [austin2016, 1.1]. Moreover, Pansu showed that asymptotic cones of a simply connected nilpotent Lie group $G$ are isomorphic to its Carnotification $\mathrm{gr}(G)$ [pansu89-carnot]. Combining these two results, we obtain that two $\mathrm{L}^{1}$-ME finitely generated nilpotent groups have the same associated Carnot. On the other hand, Delabie, Llosa Isenrich and Tessera showed a (slightly stronger) converse to this result: namely, two virtually nilpotent finitely generated groups with isomorphic Carnot are $\textrm{L}^{p}$-OE for some $p>1$ [delabie-llosa-tessera]. We sum up these results in the following statement:
@@
[delabie-llosa-tessera, 1.5] Let $\Gamma$ and $\Lambda$ be two finitely generated virtually nilpotent groups. The following are equivalent:
@@ 4.4 Examples of nilpotent groups with different Betti numbers via filiform algebras
We will describe some examples of nilpotent Lie algebras for which their Lie algebra cohomology is understood. All of the examples we will present come from [shalom2004harmonic], [garcia-llosa-pallier] and [delabie-llosa-tessera]. More examples are provided in [gotfredsen-kyed]. We could maybe enrich the list of examples by looking at the concrete bounds on filling functions provided in [young-filling-nilpotent].
@@
We call $L_{n}$ and $\mathrm{gr}(L_{n})$ the simply connected nilpotent Lie groups integrating $\mathfrak{l}_{n}$ and $\mathrm{gr}(\mathfrak{l}_{n})$. Let $\Lambda_{n}<L_{n}$ and $\Lambda_{n}^{\prime}<\mathrm{gr}(L_{n})$ be two cocompact lattices inside $L_{n}$ and $\mathrm{gr}(L_{n})$ respectively. By Mal’cev’s theorem [raghunathan, 2.18], the groups $\Lambda_{n}$ and $\Lambda_{n}^{\prime}$ are finitely generated, nilpotent and torsion-free.
The groups $L_{n}$ and $\mathrm{gr}(L_{n})$ are $\textrm{L}^{p}$-OE for some (small) $p>1$ by [delabie-llosa-tessera, 1.4] and hence $\Lambda_{n}$ and $\Lambda_{n}^{\prime}$ are also $\textrm{L}^{p}$-OE for the same $p$. On the other hand, as already pointed out in [shalom2004harmonic, p. 152], these Lie algebras satisfy
@@
[garcia-llosa-pallier, Table 2, see the groups $L_{5,7}\times_{Z}L_{3,2}$ and $L_{5,6}\times_{Z}L_{3,2}$]. Corollary 4.5 implies that the corresponding groups cannot be mutually coboundedly $\textrm{L}^{p}$-ME for large $p>1$. Notice that none of these two central products of Lie algebras is Carnot, but since they have the same graded algebras
@@
the groups $L_{5}\times_{Z}L_{3}$ and $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ must be $\textrm{L}^{p}$-OE for some $p>1$ [delabie-llosa-tessera, 1.4]. It is worth noting that the difference between $L_{5}\times_{Z}L_{3}$ and $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ is somewhat harder to detect than the difference between $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ and $\mathrm{gr}(\mathrm{gr}(L_{5})\times_{Z}L_{3})$, as in the latter pair the second Betti number differs, while in the former pair this only happens for the third Betti number. In fact it was already known that the simply connected nilpotent Lie groups $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ and $\mathrm{gr}(\mathrm{gr}(L_{5})\times_{Z}L_{3})$ are not $\textrm{L}^{p}$-OE for any $p>5$ [delabie-llosa-tessera, 1.14] using cohomology classes in degree 2 satisfying particular conditions when interpreted as central extensions.
@@ 5 Applications for non-cocompact lattices in rank 1 simple Lie groups
The article [bader-sauer] deals with unitary cohomology of semisimple groups and their lattices. One of the main breakthroughs in their work is the use of polynomial cohomology to deal with cohomology of non-cocompact lattices. Their methods do not take into account the degrees of the polynomials appearing in their cohomologies, so when dealing with lattices they only treat higher rank lattices, as these are $\textrm{L}^{p}$-integrable for every $0<p<\infty$. A non-cocompact lattice $\Lambda$ in a rank 1 simple Lie group $G$ is $\textrm{L}^{p}$-integrable for $p<\mathrm{Confdim}(\partial G)$, but not $\textrm{L}^{p}$-integrable for $p\geq\mathrm{Confdim}(\partial G)$. This means that if we want to perform induction à la Bader-Sauer, we need some ideas from Section 2.
@@
Let $\Lambda$ be a lattice in a locally compact second countable group $G$. Let $(\rho,V)$ be an isometric representation of $\Lambda$ on some separable Banach space $V$. Suppose that for some $d\in\mathbb{N}$, the higher homological filling functions $\mathrm{cFV}_{\Lambda}^{k}$ of $\Lambda$ are polynomial for $k\leq d$. Let $N_{k}=\prod_{i=2}^{k}\deg\mathrm{cFV}_{\Lambda}^{i}$ and $1\leq p<\infty$ and suppose that the lattice $\Lambda$ is $\mathrm{L}^{pN_{k}}$-integrable in $G$. Then there exist continuous surjective linear maps:
$H^{k}(G,I^{p}(\rho))\twoheadrightarrow H^{k}(\Lambda,\rho),$
@@
Let $X$ be a Borel fundamental domain for the right $\Lambda$-action on $G$. Even if we are in a slightly different setting as in Proposition 3.2, the integrability condition on $\Lambda$ ensures in the same way that the induction map
$I:C_{\mathrm{pol}\leq N_{k}}^{k}(\Lambda,V)\to L_{\mathrm{loc}}^{p}(G^{k+1},L^{p}(X,V))^{G}$
is well-defined. The inclusion $I^{p}(V)=L^{p}(X,V)$ in the Fréchet module $L_{\mathrm{loc}}^{p}(X,V)$ (with the same action) induces a natural map:
@@
We have a continuous isomorphism $L_{\mathrm{loc}}^{p}(G^{k+1},L_{\mathrm{loc}}^{p}(X,V))^{G}\simeq L_{\mathrm{loc}}^{p}(G^{k+1},V)^{\Lambda}$. Let $\alpha(g,x):G\times X\to\Gamma$ be the lattice cocycle associated to the fundamental domain $X$, where $\alpha(g,x)\in\Gamma$ is defined as the unique element in $\Gamma$ satisfying $g^{-1}x\alpha(g,x)\in X$ for $x\in X$ and $g\in G$. We can construct a chain map $f:C(\Lambda^{k+1},V)^{\Lambda}\to L_{\mathrm{loc}}^{p}(G^{*+1},V)^{\Lambda}$ by setting for $c\in C(\Lambda^{k+1},V)^{\Lambda}$
@@
In [blanc, 3.5], it is shown that the modules $L_{\mathrm{loc}}^{p}(G^{*+1},V)$ form a relatively injective resolution by continuous $\Lambda$-modules of the continuous $\Lambda$-module $V$. Hence by the fundamental theorem of relative homological algebra, the chain map $f$ admits an inverse chain homotopy equivalence. This means that we can construct a continuous chain homotopy equivalence:
@@
Hence there is a continuous surjective map $H^{k}_{\mathrm{ct}}(G,L^{p}(X,V))\to H^{k}(\Lambda,V)$. ∎
@@
Let $\Lambda$ be a non-cocompact lattice in a simple Lie group $G$ of rank 1. In [shalom2000rigidity-semisimple, 3.7], Shalom proved that $\Lambda$ is $\textrm{L}^{q}$-integrable in $G$ for every $q<\mathrm{Confdim}(\partial G)$ (though stated more clearly in the Appendix [marrakchi-dlSalle, A.5]). Hence we can use Theorem 5.1 for any $p<\mathrm{Confdim}(\partial G)/N_{k}$.

### v2

@@ 4 Applications for virtually nilpotent groups
(Gromov [25]) A finitely generated group $(\Gamma,S)$ has polynomial growth if and only if $\Gamma$ is virtually nilpotent.
@@
Let $\Gamma$ be a finitely generated nilpotent group. Then for every $d\in\mathbb{N}_{\geq 2}$, $\Gamma$ is of type $F_{d}$ and the combinatorial filling function $\mathrm{cFV}^{d}_{\Gamma}$ is polynomially bounded.
Finiteness properties of finitely generated nilpotent groups come from the fact that they are cocompact lattices in their Mal’cev completions [33]. It is hard to single out one reference for the polynomiality of filling functions of nilpotent groups. This is stated in [26, p. 56 or depending on edition p. 82]. The book [43] allows to see it in different ways. For instance, in the book’s terminology, nilpotent Lie groups are NC [43, Section 2.2.1], and if a group is NC then it is polynomially retractable [43, Theorem 7.10] and being polynomially retractable implies having the polynomial filling property [43, Definition 7.12], where the latter means that the more classical Lipschitz versions of the filling functions $\mathrm{FV}_{\Gamma}^{k}$ are polynomially bounded for every $k\in\mathbb{N}_{\geq 2}$. The Federer-Fleming deformation technique shows that $\mathrm{cFV}_{\Gamma}^{k}\sim\mathrm{FV}_{\Gamma}^{k}$ (see for instance [5, Theorem 6.6]).
The third property that we need is the higher analogue of Shalom’s property $H_{T}$ [42, p. 125]. To our knowledge, this property was first shown for nilpotent groups by Blanc [6].
@@
(Blanc [6, 10.5], [28, p. 243]) Let $G$ be a connected and simply connected nilpotent Lie group. Then for every $k\geq 1$ and for every continuous unitary representation $\pi$ without $G$-invariant vectors we have:
@@
By Shapiro’s lemma for cocompact lattices [42, 3.2.2], this property holds for torsion-free finitely generated nilpotent groups. By Shapiro’s lemma for finite index subgroups, it also holds for general finitely generated virtually nilpotent groups.
@@ 4.1 Injectivity of induction
The following statement and its consequences will depend on the number $d_{\Gamma}=\inf\{d\in\mathbb{R}_{>0},\#S_{\Gamma}(n)\lesssim n^{d}\}$, where $S_{\Gamma}(n):=\{\gamma\in\Gamma,|\gamma|_{S}=n\}$ denotes the sphere of radius $n$ for the length function $|\cdot|$. In the introduction, we stated results using the constant $d(\Gamma)=\inf\{d\in\mathbb{R}_{>0},\#B_{\Gamma}(n)\lesssim n^{d}\}$, where $B_{\Gamma}(n):=\{\gamma\in\Gamma,|\gamma|_{S}\leq n\}$. These two differ at most by 1 as we have $d(\Gamma)-1\leq d_{\Gamma}\leq d(\Gamma)$. Conjecturally we have $d_{\Gamma}=d(\Gamma)-1$ [12, 8.2] (see also [12, 8.3]).
@@ 4.2 Betti numbers of nilpotent groups and $\textrm{L}^{p}$-ME
The next Corollary is obtained by specializing the previous Theorem to the case of the trivial representation. Its proof is essentially the same as [42, 4.1.1].
@@
Let $\Lambda$ and $\Gamma$ be finitely generated virtually nilpotent groups with the same associated Carnot group. Let $d$ be the degree of polynomial growth of both $\Lambda$ and $\Gamma$. Let $N=\max\{\prod_{i=2}^{k}\deg\mathrm{cFV_{\Lambda}^{i}},\prod_{i=2}^{k}\deg\mathrm{cFV_{\Gamma}^{i}}\}$ and let $M>\alpha_{k-1}(N)$ where the functions $\alpha_{i}$ are defined in Theorem C. Since the groups $\Lambda$ and $\Gamma$ are not hyperbolic, we can choose $\alpha_{k-1}(N)$ to be $N^{2}+(k-2)N$ when $k\geq 3$ (Corollary 2.10) and $\alpha_{k-1}(N)=\alpha_{1}(N)=N+1$ when $k=2$ (Corollary 2.8). Suppose that $\Lambda$ and $\Gamma$ admit a mutually cobounded $\textrm{L}^{p}$-ME coupling $(\Omega,X_{\Lambda},X_{\Gamma},m)$ for some $p>dk+2M+1$.
@@ 4.3 Nilpotent Lie algebras and gradings
Every finitely generated virtually nilpotent group contains a finite index subgroup that is torsion-free and nilpotent [39, Lemma 4.6]. Moreover, every finitely generated, torsion-free, nilpotent group $\Gamma$ arises as a cocompact lattice in a simply connected nilpotent Lie group $G$, called the Mal’cev completion of $\Gamma$, whose corresponding Lie algebra $\mathfrak{g}$ has a basis with rational structural constants [33] [39, 2.18]. One can compute Betti numbers of $\Gamma$ using Betti numbers of the Lie algebra $\mathfrak{g}$ (which are defined using Lie algebra cohomology). Indeed, since $\Gamma$ is cocompact in $G$, we have that
@@
for every $n$ (this is can be seen using [42, 3.2.2] and [28, p. 243]). Moreover, we also have $b_{n}(G)=b_{n}(\mathfrak{g})$ [28, Chapitre II], where the latter means Betti numbers in the sense of Lie algebra cohomology.
@@
We recall some results on $\textrm{L}^{p}$-measure equivalence of nilpotent groups. On one hand, Austin showed that $\mathrm{L}^{1}$-ME finitely generated groups have bilipschitz asymptotic cones [3, 1.1]. Moreover, Pansu showed that asymptotic cones of a simply connected nilpotent Lie group $G$ are isomorphic to its Carnotification $\mathrm{gr}(G)$ [36]. Combining these two results, we obtain that two $\mathrm{L}^{1}$-ME finitely generated nilpotent groups have the same associated Carnot. On the other hand, Delabie, Llosa Isenrich and Tessera showed a (slightly stronger) converse to this result: namely, two virtually nilpotent finitely generated groups with isomorphic Carnot are $\textrm{L}^{p}$-OE for some $p>1$ [17]. We sum up these results in the following statement:
@@
[17, 1.5] Let $\Gamma$ and $\Lambda$ be two finitely generated virtually nilpotent groups. The following are equivalent:
@@ 4.4 Examples of nilpotent groups with different Betti numbers via filiform algebras
We will describe some examples of nilpotent Lie algebras for which their Lie algebra cohomology is understood. All of the examples we will present come from [42], [23] and [17]. More examples are provided in [24]. We could maybe enrich the list of examples by looking at the concrete bounds on filling functions provided in [47].
@@
We call $L_{n}$ and $\mathrm{gr}(L_{n})$ the simply connected nilpotent Lie groups integrating $\mathfrak{l}_{n}$ and $\mathrm{gr}(\mathfrak{l}_{n})$. Let $\Lambda_{n}<L_{n}$ and $\Lambda_{n}^{\prime}<\mathrm{gr}(L_{n})$ be two cocompact lattices inside $L_{n}$ and $\mathrm{gr}(L_{n})$ respectively. By Mal’cev’s theorem [39, 2.18], the groups $\Lambda_{n}$ and $\Lambda_{n}^{\prime}$ are finitely generated, nilpotent and torsion-free.
The groups $L_{n}$ and $\mathrm{gr}(L_{n})$ are $\textrm{L}^{p}$-OE for some (small) $p>1$ by [17, 1.4] and hence $\Lambda_{n}$ and $\Lambda_{n}^{\prime}$ are also $\textrm{L}^{p}$-OE for the same $p$. On the other hand, as already pointed out in [42, p. 152], these Lie algebras satisfy
@@
[23, Table 2, see the groups $L_{5,7}\times_{Z}L_{3,2}$ and $L_{5,6}\times_{Z}L_{3,2}$]. Corollary 4.5 implies that the corresponding groups cannot admit mutually cobounded $\textrm{L}^{p}$-ME couplings for large $p>1$. Notice that none of these two central products of Lie algebras is Carnot, but since they have the same graded algebras
@@
the groups $L_{5}\times_{Z}L_{3}$ and $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ must be $\textrm{L}^{p}$-OE for some $p>1$ [17, 1.4]. It is worth noting that the difference between $L_{5}\times_{Z}L_{3}$ and $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ is somewhat harder to detect than the difference between $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ and $\mathrm{gr}(\mathrm{gr}(L_{5})\times_{Z}L_{3})$, as in the latter pair the second Betti number differs, while in the former pair this only happens for the third Betti number. In fact it was already known that the simply connected nilpotent Lie groups $\mathrm{gr}(L_{5})\times_{Z}L_{3}$ and $\mathrm{gr}(\mathrm{gr}(L_{5})\times_{Z}L_{3})$ are not $\textrm{L}^{p}$-OE for any $p>5$ [17, 1.14] using cohomology classes in degree 2 satisfying particular conditions when interpreted as central extensions.
@@ 5 Applications for non-cocompact lattices in rank 1 simple Lie groups
The article [5] deals with unitary cohomology of semisimple groups and their lattices. One of the main breakthroughs in their work is the use of polynomial cohomology to deal with cohomology of non-cocompact lattices. Their methods do not take into account the degrees of the polynomials appearing in their cohomologies, so when dealing with lattices they only treat higher rank lattices, as these are $\textrm{L}^{p}$-integrable for every $0<p<\infty$. A non-cocompact lattice $\Lambda$ in a rank 1 simple Lie group $G$ is $\textrm{L}^{p}$-integrable for $p<\mathrm{Confdim}(\partial G)$, but not $\textrm{L}^{p}$-integrable for $p\geq\mathrm{Confdim}(\partial G)$. This means that if we want to perform induction as in [5, Section 6], we need some ideas from Section 2.
@@
Let $\Lambda$ be a lattice in a locally compact second countable group $G$. Let $(\rho,V)$ be an isometric representation of $\Lambda$ on some separable Banach space $V$. Suppose that for some $d\in\mathbb{N}$, the higher homological filling functions $\mathrm{cFV}_{\Lambda}^{k}$ of $\Lambda$ are polynomially bounded for $k\leq d$. Let $N_{k}=\prod_{i=2}^{k}\deg\mathrm{cFV}_{\Lambda}^{i}$ and $1\leq p<\infty$ and suppose that the lattice $\Lambda$ is $\mathrm{L}^{pN_{k}}$-integrable in $G$. Then there exist continuous surjective linear maps:
$H_{\mathrm{ct}}^{k}(G,I^{p}(\rho))\twoheadrightarrow H^{k}(\Lambda,\rho),$
$\overline{H}_{\mathrm{ct}}^{k}(G,I^{p}(\rho))\twoheadrightarrow\overline{H}^{k}(\Lambda,\rho).$
@@
Our proof uses the same setting as in [5, Section 6.5]. Let $X$ be a Borel fundamental domain for the right $\Lambda$-action on $G$ and let $\alpha:G\times X\to\Gamma$ be its corresponding lattice cocycle (with the convention that $g^{-1}x\alpha(g,x)\in X$ for $x\in X$ and $g\in G$). We define induction as in [5, 6.23]: for $c\in C^{k}(\Lambda,V)^{\Lambda}$ and $g_{0},\ldots,g_{k}\in G$, let $I(c)(g_{0},\ldots,g_{k}):X\to V$ be the measurable function defined by
$I(c)(g_{0},\ldots,g_{k})(x)=c(\alpha(g_{0},x),\ldots,\alpha(g_{k},x)).$
Even if we are in a slightly different setting from Proposition 3.2, the integrability condition on $\Lambda$ ensures in the same way that the induction map
$I:C_{\mathrm{pol}\leq N_{k}}^{k}(\Lambda,V)^{\Lambda}\to L_{\mathrm{loc}}^{p}(G^{k+1},L^{p}(X,V))^{G}$
is well-defined. The inclusion $I^{p}(V)=L^{p}(X,V)$ into the Fréchet module $L_{\mathrm{loc}}^{p}(X,V)$ (with the same action) induces a natural map:
@@
We have a continuous isomorphism $L_{\mathrm{loc}}^{p}(G^{k+1},L_{\mathrm{loc}}^{p}(X,V))^{G}\simeq L_{\mathrm{loc}}^{p}(G^{k+1},V)^{\Lambda}$. We can construct a chain map $f:C(\Lambda^{k+1},V)^{\Lambda}\to L_{\mathrm{loc}}^{p}(G^{*+1},V)^{\Lambda}$ by setting for $c\in C(\Lambda^{k+1},V)^{\Lambda}$
@@
In [6, 3.5], it is shown that the modules $L_{\mathrm{loc}}^{p}(G^{*+1},V)$ form a relatively injective resolution by continuous $\Lambda$-modules of the continuous $\Lambda$-module $V$. Hence, by the fundamental theorem of relative homological algebra, the chain map $f$ admits an inverse chain homotopy equivalence. This means that we can construct a continuous chain homotopy equivalence:
@@
Hence there is a continuous surjective map $H^{k}_{\mathrm{ct}}(G,L^{p}(X,V))\to H^{k}(\Lambda,V)$.
By continuity of $I$ and $T$, we have the same commutative diagram for reduced cohomology. The comparison map $\overline{H}_{\mathrm{pol}:N_{k}\to N_{k}}^{k}(\Lambda,V)\to\overline{H}^{k}(\Lambda,V)$ is also surjective by Theorem 2.1, hence we obtain a continuous surjective map $\overline{H}^{k}_{\mathrm{ct}}(G,L^{p}(X,V))\to\overline{H}^{k}(\Lambda,V)$. ∎
@@
Let $\Lambda$ be a non-cocompact lattice in a simple Lie group $G$ of rank 1. In [41, 3.7], Shalom proved that $\Lambda$ is $\textrm{L}^{q}$-integrable in $G$ for every $q<\mathrm{Confdim}(\partial G)$ (though stated more clearly in the Appendix [34, A.5]). Hence we can use Theorem 5.1 for any $p<\mathrm{Confdim}(\partial G)/N_{k}$.